# sample.bib

### v1

@@ 2.1 Human-like Driving Based on Language and Vision Models
With the significant research advances in language models T5[42], GPT-4[43], Llama[4], vision models CLIP[44], ALIGN[45], Segment anything[46], and multimodal models BLIP-2[47], LLaVA[48], and GPT-4V[49] as well as other FMs, its powerful reasoning capability is considered to have brought a new dawn for realizing Artificial General Intelligence (AGI)[50], which has had a significant and far-reaching impact on all aspects of society. In autonomous driving, FMs such as language and vision also show great potential, which is expected to enhance the understanding and reasoning of autonomous driving models on driving scenarios and realize human-like driving for autonomous driving.
@@
In the following, we will provide a review and introduction of the research related to the enhancement of the understanding of the driving scenarios by the autonomous driving system based on FMs of language and vision, as well as the reasoning to give language-guided instructions and driving actions, as illustrated in Fig. 3. Related work on enhancing the understanding of driving scenes is presented in Section 2.1.1, on reasoning to give verbal commands in 2.1.2, and on reasoning to give action commands in 2.1.3.
@@
Figure 3: The pipeline diagram for enhancing autonomous driving leveraging FMs, where FMs refer to language models and vision models. FMs can learn the perceptual information and utilize their powerful ability to understand the driving scenarios and reason to give language-guided instructions and driving actions to enhance autonomous driving.
2.1.1 Understanding of Driving Scenarios
@@
The study by Vasudeva et al.[51] found that the model’s understanding of the scene and object localization can be effectively enhanced by acquiring verbal descriptions and gaze estimation, etc. Li et al.[52] proposed an image captioning model that generates high-level semantic information to enhance the model’s understanding of the traffic scene. Their work verified that linguistic and visual features can effectively enhance the model’s understanding of driving scenes.
@@
Chen et al.[53] transferred image and text features into a 3D point cloud network based on CLIP, which enhanced the model’s understanding of the 3D scenes. Sriram et al.[54] designed and proposed an autonomous navigation framework through a combination of local semantics and natural language commands by fleveraging the natural language coding with the coding of 3D semantic maps obtained from visual images, which was experimentally validated in the CARLA simulator and KITTI dataset[55] for experimental validation. Romero et al.[56] built a video analytics system leveraging a visual-linguistic model, which effectively improves query accuracy based on the powerful comprehension of the visual-linguistic model. Elhafsi et al.[57] improved query accuracy by converting observed visual information into natural language descriptions and passing them to a large language model to identify semantic anomalies leveraging its powerful reasoning ability. Dewangan et al.[58] proposed a method to enhance BEV maps by detecting the features of each object in the BEV through a visual-linguistic multimodal FM (Blip-2[47], Minigpt-4[59], Instructblip[60]) and through linguistic characterizations to obtain a language-enhanced BEV map.
2.1.2 Language-Guided Instructions
Here we present a review of studies that give linguistic commands through FMs, mainly descriptive commands, such as “Red light ahead, you should slow down”, “Intersection ahead, please pay attention to pedestrians”, etc.
@@ 2.1.2 Language-Guided Instructions
Ding et al.[61] used a visual coder to encode video data, which was then fed into a large language model to generate corresponding driving scenario descriptions and suggestions. In particular, this work also proposed a method that enables high-resolution feature maps and the obtained high-resolution information to be fused into multimodal large language models (M-LLMs) to further enhance the model’s recognition, interpretation, and localization capabilities. Fu et al.[62] explored the potential of leveraging LLMs to understand driving environments in a human-like manner, utilizing the LLaMA-Adapter[63] to describe the scene data, and then giving linguistic commands via GPT-3.5. Wen et al.[64] proposed DiLu, a knowledge-driven paradigm based on previous work that can make decisions based on common-sense knowledge and accumulate experience. In particular, the article pointed out that DiLu possesses the ability to direct experience acquisition of real-world data, which has the potential for practical deployment of autonomous driving systems. In the industry, Wayve proposed LINGO-1[65], a grand model of self-driving interaction based on a visual-verbal-action grand model, where the model can interpret itself and answer visually while driving. It introduced human driving experiences, which can explain various causal elements in the driving scenarios through natural language descriptions, acquire feature information in the driving scenario in a human-like understanding way, and learn and give interactive language commands.
2.1.3 Generation of Actions
@@
As described in the previous section 2.1.2, academia and industry have attempted to embed GPT linguistic knowledge into autonomous driving decisions to improve the performance of autonomous driving in the form of linguistic commands to promote the application of large models to autonomous driving. Long before FM made breakthroughs in the LLM field, some works attempted to improve the performance of autonomous driving through similar research ideas. For example, the MP3 framework proposed by Casas et al.[66] used high-level semantic information as a decision training guide, which together with sensory data constitutes the input to build algorithms to realize motion prediction.
@@
The research on the application of LLMs in autonomous driving is on the ascendant, and the GPT series, as the most successful variant of the transformer architecture, may be able to bring breakthroughs to improve the comprehensive performance at multiple levels. LLM is a representative of FMs from the level of linguistic knowledge that empowers the development of autonomous driving, however, linguistic descriptions and reasoning are not directly applied by the autonomous driving system. Considering that the large model is expected to be truly deployed at the vehicle end ,it needs to eventually fall on the planning or control instructions, i.e., FMs should eventually empower the autonomous driving from the action state level. However, how to quantize linguistic decisions into action commands, such as planning and control, available to the autonomous driving system still faces great challenges. Some scholars have already made preliminary explorations, but there is still much room for development. What’s more, some scholars have explored the construction of autonomous driving models through a GPT-like approach, which directly outputs trajectories and even control commands based on LLM. In Table LABEL:tab:1 we provide a brief overview of some representative works.
@@
Sha et al.[67] proposed LanguageMPC, which employs GPT-3.5 as a decision-making component for complex autonomous driving scenarios that require human common sense understanding. By designing cognitive pathways for integrated reasoning in LLM, sha et al. proposed algorithms to transform LLM decisions into actionable driving control commands, which improved the vehicle’s ability to handle complex driving behaviors. Jain et al.[68] achieved navigation localization and further planning of trajectories with the help of visual perception for explicit verbal commands. Omama et al.[69] constructed a multimodal map-based navigation and localization method called ALT-Pilot, which can be used to navigate to arbitrary destinations without the need for high-definition LiDAR maps, demonstrating that off-the-shelf visual language models can be used to construct linguistically enhanced terrain maps.
Some scholars have also attempted to construct autonomous driving models directly through a GPT-like approach, i.e., leveraging LLMs to construct an end-to-end autonomous driving planner, which directly outputs predicted trajectories, path planning, and even control commands, intending to effectively improve the ability of autonomous driving models to generalize to unknown driving scenarios. Keysan et al.[70] took information about the current surroundings on the road as input and then output a sequence that includes future vehicle decision instructions and spatial paths that can be followed by the self-driving vehicle. Seff et al.[71] proposed MotionLM, which uses motion prediction as a language modeling task to learn multimodal distributions by representing continuous trajectories as discrete sequences of motion tokens leveraging a single standard language modeling objective to predict the future behaviors of road network participants. Mao et al.[72] proposed the GPT-Driver model to reformulate motion planning as a language modeling problem by representing the inputs and outputs of the planner as linguistic tokens and leveraging the LLM to generate driving trajectories through linguistic descriptions of coordinate positions. Furthermore, the model setup allows for the description of internal decision-making processes in natural language, which also improves the interpretability of the decisions. Pallagani et al.[73] constructed Plansformer, which is both a large language model and a planner, showing the great potential of a large language model fine-tuned as a planner from a variety of planning tasks. Wang et al.[74] constructed the BEVGPT model, which takes as input information about the current surroundings on the road and then outputs a sequence that includes future vehicle decision instructions and spatial paths that can be followed by self-driving vehicles. DriveGPT4 proposed by Xu et al.[75] is further used for autonomous driving planning control by inputting video data, textual data, and historical control signals and outputs as textual responses to the problem as well as predictions of control signals to be used for vehicle operation. Cui et al.[76] utilized GPT-4 with inputs of natural language descriptions and environment perception data to make LLM directly output driving decisions and operation commands, and further experimented with highway overtaking and lane changing scenarios in Ref.[77] to compare driving decisions provided by LLM with different cues, and the study showed that chained-thinking cueing helps LLM to make better driving decisions. Considering the scale challenges of Visual Language Model (VLM), Chen et al.[78], based on the idea that digital vector modalities are more compact than image data, fused vectorized 2D scene representations with pre-trained LLMs to improve the LLM’s ability to interpret and reason about the integrated driving situation, giving scene interpretation and vehicle control commands. Mao et al.[mao2023language] proposed Agent Driver, which utilized LLM to introduce a general-purpose library of tools accessible via function calls, cognitive memory for common sense and empirical knowledge for decision making, and a reasoning machine capable of chain-of-thinking (CoT) reasoning, task planning, motion planning, and self-reflection to achieve a more nuanced, human-like approach to autopilot.
@@
This type of research idea is much closer to human driving than pure knowledge embedding to make an autonomous driving model and perhaps has the potential to become one of the main development directions in the future with the development of large models. However, it should be noted that these new frameworks are also questionable in terms of reliability due to the unresolved pitfalls of large models themselves, such as “illusions” (LLMs may generate content that conflicts with source or factual information). Specific details about the problems of the large models themselves and the challenges inherited in autonomous driving will be discussed in detail in Section 5.
@@
Omama et al.[69] 2023 OSM Maps with Descriptions, LiDAR, Camera Location, Trajectory SL This work utilizes a Bayesian state estimation model fleveraging visual-linguistic features to generate global paths, plan trajectories, and control vehicles to complete navigation.
Keysan et al.[70] 2023 Scene Raster, Text Prompt Trajectory SL This work encodes the rasterized scene and the text prompt with pre-trained models dedicated for each modality, finally picks the target trajectory from a pre-generated trajectory set.
Seff et al.[71] 2023 Multimodal Scene Motion Prediction SL This work uses a single standard linguistic modeling objective to learn multimodal distributions for predicting the future behavior of traffic participants.
@@
Wang et al.[74] 2023 Scene Information Driving Decisions, Trajectory SL In pre-training stage, this work trains a causal transformer for driving scenario prediction and decision-making. In fine-tuning stage, it adapts to motion planning and accurate BEV generation.

### v2

@@ 3 Human-like Driving Based on Language and Vision Models
With significant research progress in LLMs BERT, GPT-4, Llama[44], Vision Language Models (VLMs) CLIP[4], ALIGN[45], BLIP-2[47] and Multi-modal Large Language Models(M-LLMs) GPT-4V[49], LLaVA[48] and Gemini[team2023gemini], as well as other FMs, their powerful reasoning capabilities are considered to have ushered in a new dawn for the realization of Artificial General Intelligence (AGI)[50], which has had a significant and far-reaching impact on all aspects of society. In autonomous driving, FMs such as language and vision also show great potential, which is expected to improve the understanding and reasoning ability of autonomous driving models on driving scenarios and realize human-like driving for autonomous driving.
@@
We provide an introduction of the research related to the enhancement of the understanding of the driving scenarios by the autonomous driving system based on FMs of language and vision, as well as the reasoning to give language-guided instructions and driving actions, as illustrated in Fig. 4. Related work on enhancing the understanding of driving scenarios is presented in Section 3.1, on reasoning to give language-guided instructions in Section 3.2, and on reasoning to generate driving actions in Section 3.3.
@@
Figure 4: The pipeline diagram for enhancing autonomous driving leveraging FMs, where FMs refer to language models and vision models. FMs can learn perceptual information and utilize their powerful ability to understand the driving scenarios and reason to give language-guided instructions and driving actions to enhance autonomous driving.
3.1 Understanding of Driving Scenarios
@@
The study by Vasudeva et al.[51] found that the ability of the model to comprehend the scene and localize objects can be effectively enhanced by acquiring verbal descriptions and gaze estimation, etc. Li et al.[52] proposed an image captioning model that generates high-level semantic information to improve its comprehension of the traffic scene. Their work verified that linguistic and visual features can effectively enhance the comprehension of driving scenarios.
@@
Sriram et al.[54] have proposed an autonomous navigation framework that combines semantic segmentation results with natural language commands. This framework has been verified to be effective as a vehicle driver in the CARLA simulator and the KITTI dataset[126]. Elhafsi et al.[57] identified semantic anomalies by converting observed visual information into natural language descriptions and passing them to the LLM to exploit its powerful reasoning capabilities. In the context of VLMs applications, Chen et al.[53] transferred image and text features to a 3D point cloud network based on CLIP to enhance the model’s understanding of the 3D scene. Romero et al.[56] constructed a video analytics system based on VIVA[romero2022optimizing], an extended model of CLIP, intending to improve query accuracy through the utilization of the powerful comprehension of VLM. Tian et al.[tian2024drivevlm] employed VLM to describe and analyze the driving scenarios, thereby enhancing the understanding of the driving scenarios. In addition to direct enhancement of scene data, perceptual features have also been explored for enhancement. Pan et al.[pan2024vlp] designed the Ego-car prompt to enhance the obtained BEV features using the language model in CLIP. Dewangan et al.[58] proposed an approach to enhance BEV maps by detecting the features of each object in the BEV through VLMs (Blip-2[47], Minigpt-4[59], Instructblip[60]) and through linguistic characterizations to obtain a language-enhanced BEV map. However, existing VLMs are constrained to the 2D domain, lacking the capacity for spatial awareness and long-horizon extrapolation. To address this issue, Zhou et al.[zhou2024embodied] proposed a model——Embodied Language Model (ELM), which enhances the understanding of driving scenarios over long-time domains and across space. This is achieved by using diverse pre-training data and selecting adaptive tokens.
3.2 Language-Guided Instructions
@@
Here we present a review of studies that give linguistic instructions through FMs, mainly descriptive instructions, such as “Red light ahead, you should slow down”, “Intersection ahead, please pay attention to pedestrians”, etc. Ding et al.[61] used a visual coder to encode video data, which was then fed into a large language model to generate corresponding driving scenario descriptions and suggestions. In particular, this work also proposed a method that enables high-resolution feature maps and the obtained high-resolution information to be fused into M-LLMs to further enhance the model’s recognition, interpretation, and localization capabilities. Fu et al.[62] explored the potential of leveraging LLMs to comprehend driving environments like a human being, utilizing the LLaMA-Adapter[63] to describe the scene data, and then giving linguistic commands via GPT-3.5. Wen et al.[64] proposed DiLu, a knowledge-driven paradigm based on previous work that can make decisions based on common-sense knowledge and accumulate experience. In particular, the article pointed out that DiLu possesses the ability to direct experience acquisition of real-world data, which has the potential for practical deployment of autonomous driving systems. To further improve the safety of LLM-based autonomous driving, Wang et al.[wang2023empowering] used an MPC-based verifier to evaluate and provide feedback on trajectory planning, and then fused prompt learning to enable LLM to perform in-context safety learning, which overall improved the safety and reliability of autonomous driving. In order to enrich the data input to obtain more accurate scene information, Wang et al.[wang2023drivemlm] utilized multi-model LLM to enable the autonomous driving system to obtain linguistic commands. Meanwhile, for the gap between linguistic commands and vehicle control commands, this work performed an alignment operation on decision states.
The aforementioned works are more in the context of data sets and simulation environments, and there has been some exploratory work in terms of real vehicle testing. Wayve proposed LINGO-1[65], a grand model of self-driving interaction based on a visual-verbal-action grand model, where the model can interpret itself and answer visually while driving. It introduced human driving experiences, which can explain various causal elements in the driving scenarios through natural language descriptions, acquire feature information in the driving scenario in a human-like understanding way, and learn and give interactive language commands. Cui et al.[cui2023large] innovatively placed the LLM in the cloud, entered human commands, and leveraged the reasoning ability of the LLM to generate executable code instructions. However, the work suffers from latency issues and has room for improvement in terms of real-time performance requirements for autonomous driving.
The pipeline for incorporating LLMs into autonomous driving systems in more current research is summarized in general terms in Fig.4, which is mainly implemented through scene understanding, high-level semantic decision-making, and trajectory planning. In this subsection, we summarize the advanced decision-making applications and argue that the research processes have some similarities. To more clearly illustrate how they work, we use DriveMLM[wang2023drivemlm], a typical recent research work, as an example for further illustration in Fig.5.
Figure 5: For the application of LLMs to autonomous driving system decision-making, a typical pipeline is shown in this figure, referenced from DriveMLM[wang2023drivemlm].
DriveMLM simulates the behavioral planning module of a modular autopilot system by using a Multi-modal LLM (M-LLM), which performs closed-loop autonomous driving in a realistic simulator based on processed perceptual information and command requirements. DriveMLM also generates natural language explanations of its driving decisions, thereby increasing the transparency and trustworthiness of the system.
3.3 Generation of Actions
@@
As described in the previous Section 3.2, academia and industry have attempted to embed GPT linguistic knowledge into autonomous driving decisions to enhance the performance of autonomous driving in the form of linguistic instructions to promote the application of FMs to autonomous driving. Long before FMs made breakthroughs in the LLMs field, some works attempted to improve the performance of autonomous driving through similar research ideas. For example, the MP3 framework proposed by Casas et al.[66] used high-level semantic information as a decision training guide, which together with sensory data constitutes the input to build algorithms to realize motion prediction.
@@
The research on the application of LLMs in autonomous driving is on the ascendant, and the GPT series, as the most successful variant of the transformer architecture, may be able to bring breakthroughs to improve the comprehensive performance at multiple levels. LLM is a representative of FMs from the level of linguistic knowledge that empowers the development of autonomous driving, however, linguistic descriptions and reasoning are not directly applied by the autonomous driving system. Considering that the large model is expected to be truly deployed at the vehicle end, it needs to eventually fall on the planning or control instructions, i.e., FMs should eventually empower autonomous driving from the action state level. Nevertheless, how to quantize linguistic decisions into action commands, such as planning and control, available to the autonomous driving system still faces great challenges. Some scholars have already made preliminary explorations, but there is still much room for development. What’s more, some scholars have explored the construction of autonomous driving models through a GPT-like approach, which directly outputs trajectories and even control commands based on LLM. In Table LABEL:tab:1 we provide a brief overview of some representative works.
@@
Sha et al.[67] proposed LanguageMPC, which employs GPT-3.5 as a decision-making module for complex autonomous driving scenarios that require human common sense comprehension. By designing cognitive pathways for integrated reasoning in LLM, sha et al. proposed algorithms to transform LLM decisions into actionable driving control commands, which improved the vehicle’s ability to handle complex driving behaviors. Jain et al.[68] achieved navigation localization and further planning of trajectories with the help of visual perception for explicit verbal commands. Omama et al.[69] constructed a multi-modal map-based navigation and localization method called ALT-Pilot, which can be used to navigate to arbitrary destinations without the need for high-definition LiDAR maps, demonstrating that off-the-shelf visual language models can be used to construct linguistically enhanced terrain maps. Pan et al.[pan2024vlp] proposed the VLP method to improve the contextual reasoning for visual perception and motion planning of an autonomous driving system with the powerful reasoning capability of LLM in the training phase, and achieved excellent performance in the open-loop end-to-end motion planning task.
Some scholars have also attempted to construct autonomous driving models directly through a GPT-like approach, i.e., leveraging LLMs to construct an end-to-end autonomous driving planner, which directly outputs predicted trajectories, path planning, and even control commands, intending to effectively improve the ability of autonomous driving models to generalize to unknown driving scenarios.
Pallagani et al.[73] constructed Plansformer, which is both a large language model and a planner, showing the great potential of a large language model fine-tuned as a planner from a variety of planning tasks. Wang et al.[74] constructed the BEVGPT model, which takes as input information about the current surroundings on the road and then outputs a sequence that includes future vehicle decision instructions and spatial paths that can be followed by self-driving vehicles.
Some works[70, drivelm, 75, 76, 77, lmdrive] took both text prompts and information about the current surroundings on the road as inputs and then output textual responses or interpretations and a sequence that includes future vehicle decision instructions and spatial paths that can be followed by the self-driving vehicle. Among them, Cui et al.[76] utilized GPT-4 with inputs of natural language descriptions and environment perception data to make LLM directly output driving decisions and operation commands, and further experimented with highway overtaking and lane changing scenarios in Ref.[77] to compare driving decisions provided by LLM with different cues, and the study showed that chained-thinking cueing helps LLM to make better driving decisions.
Some scholars have also tried different ideas. Seff et al.[71] proposed MotionLM, which uses motion prediction as a language modeling task to learn multi-modal distributions by representing continuous trajectories as discrete sequences of motion tokens leveraging a single standard language modeling objective to predict the future behaviors of road network participants. Mao et al.[72] proposed the GPT-Driver model to reformulate the motion-planning task as a language modeling problem by representing the inputs and outputs of the planner as linguistic tokens and leveraging the LLM to generate driving trajectories through linguistic descriptions of coordinate positions. Furthermore, they [mao2023language] proposed Agent Driver, which utilized LLM to introduce a general-purpose library of tools accessible via function calls, cognitive memory for common sense and empirical knowledge for decision-making, and a reasoning machine capable of Chain-of-Thought (CoT) reasoning, task planning, motion planning, and self-reflection to achieve a more nuanced, human-like approach to autopilot. Ma et al.[ma2023dolphins] proposed Dolphins. It is capable of performing tasks such as scene understanding, behavior prediction, and trajectory planning. This work demonstrates the ability of a visual language model to provide a comprehensive understanding of complex and open-world long-tail driving scenarios and solve a range of AV tasks, as well as emergent human-like capabilities including context-learning gradient-free immediate adaptation and reflective error recovery.
Considering the scale challenges of the Visual Language Model (VLM), Chen et al.[132], based on the idea that digital vector modalities are more compact than image data, fused vectorized 2D scene representations with pre-trained LLMs to improve the LLM’s ability to interpret and reason about the integrated driving situation, giving scene interpretation and vehicle control commands. Tian et al. [tian2024drivevlm] propose DriveVLM, which, through the CoT mechanism, is not only able to generate descriptions and analyses of the scenes presented in image sequences to make driving decision guidance but also further enables trajectory planning in conjunction with the traditional automated driving pipeline. The proposed work also provides possible solutions to the challenges inherent in VLM in terms of spatial reasoning and computation, realizing an effective transition between existing autopilot approaches and large model-based approaches.
As in the previous subsection, for the research work on the application of LLMs to the direct generation of trajectory planning for autonomous driving systems, we take the example of a typical recent research work in Fig.6, LMDrive[lmdrive], to hopefully illustrate more clearly how it works. LMDrive is based on the Carla simulator, and the model training consists of two phases: pre-training and command fine-tuning. In the pre-training phase, prediction headers are added to the vision encoder to perform pre-training tasks. After the pre-training is completed, the prediction headers are discarded and the vision encoder is frozen. In the instruction fine-tuning stage, the navigation instruction and the notice instruction are configured for each driving segment, and the visual tokens are processed through the time series of instruction encoding by LLaMA, and together with the textual tokens, they are inputted into LLM to obtain the prediction tokens. After the 2-MLP Adapter, the output is the planning of the future trajectory of the auto-vehicle and the flag of whether the instruction is completed or not, and the planned trajectory completes the closed-loop simulation through the transverse and longitudinal PID controllers.
Figure 6: For the application of LLMs to autonomous driving system planning, a typical pipeline is shown in this figure, referenced from LMDrive[lmdrive].
@@
This type of research idea is much closer to human driving than pure knowledge embedding to make an autonomous driving model. With the development of large models, it perhaps has the potential to become one of the main development directions in the future. Motion planning, as one of the fundamental topics in the field of intelligent robotics[wulker2019quantizing], is significant to quantifying linguistic decisions into action commands such as planning and even control available for autonomous driving systems through LLM. However, it should be noted that these new frameworks are also questionable in terms of reliability due to the unresolved pitfalls of large models themselves, such as “illusions” (LLMs may generate content that conflicts with source or factual information). Specific details about the problems of the large models themselves and the challenges inherited in autonomous driving will be discussed in detail in Section 6.
@@
Omama et al.[69] 2023 OSM Maps with Descriptions, LiDAR, Camera Location, Trajectory SL This work utilizes a Bayesian state estimation model leveraging visual-linguistic features to generate global paths, plan trajectories, and control vehicles to complete navigation.
Keysan et al.[70] 2023 Scene Raster, Text Prompt Trajectory SL This work encodes the driving scene and text prompt with pre-trained models dedicated for each modality, finally sifts through the set of trajectories to find the target trajectory.
Seff et al.[71] 2023 Multi-modal Scene Motion Prediction SL This work uses a single standard linguistic modeling objective to learn multi-modal distributions for predicting the future behaviors of traffic participants.
@@
Wang et al.[74] 2023 Scene Information Driving Decisions, Trajectory SL In the pre-training stage, this work trains a causal transformer for driving scenario prediction and decision-making. In the fine-tuning stage, it adapts to motion planning and accurate BEV generation.